# Structural Asymmetry and Rigidity of Bounded Solutions for a Parametrized Complex Linear Differential Equation

### v3

@@ Abstract
We introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter within the critical strip. We establish a rigidity phenomenon: Two solutions with the same initial condition equal to 11, corresponding to two different parameters lying on a horizontal line in the critical strip, cannot both be bounded on $1,+\infty)$. In particular, when $\Re(w)\neq\tfrac{1}{2}$, the two solutions corresponding respectively to the parameters $w$ and $1-\bar{w}$, both taken with initial condition 11, cannot be simultaneously bounded. This result provides an answer to the Dynamical Conjecture formulated in [2].
@@ 1 Main results
In this manuscript, we study the following non-homogeneous linear complex parametrized differential equation
$\dot{x}=w\,t^{-1}x+(1-w)\,t^{-1}\eta(t),$ (1)
@@
where the parameter $w\in\mathbb{C}$. The the nonhomogeneous term $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function satisfying the following hypothesis (H):
$|\eta(t)|<1,\quad\forall t\geq 1,\quad\text{and}\quad\exists\rho\in\mathbb{C}^{*}:\ \sup_{t\geq 1}\left|\int_{1}^{t}(\eta(u)-\rho)\,du\right|<+\infty.$ (H)
We call $\rho$ the rotation number of the function $\eta$. The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions. More precisely, our aim is to prove that for every $w,w^{\prime}\in\mathbb{C}$ with $\Re(w)\neq\Re(w)$ and $\Im(w)=\Im(w^{\prime})$, the solution $x_{w}$ and the solution $x_{w^{\prime}}$ with initial conditions $x_{w}(1)=x_{1-\overline{w}}(1)=1$ cannot both be bounded on $1,+\infty)$. In the study that follows, we fix a specific function $\eta$ which is locally integrable satisfying hypothesis (H) with a fixed rotation number $\rho\in\mathbb{C}^{*}$. The function
$t\mapsto t^{w}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
@@
$\psi_{w}(z,t)=t^{w}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
The analysis in this manuscript is valid only when the real part of the parameter is in $(0,1)$. We consider the following parameter set,
$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)\in(0,1),\quad\Im(w)\neq 0\Big\}.$
@@
For every $w\in B$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1). The point $z$ is called the initial condition of the solution $\psi_{w}(z,\cdot)$.
@@
We denote by $\mu$ the function $\mu:B\mapsto\mathbb{C}$ defined as
@@
where the function $\mu$ is given by the Notation 2. Further,
$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
@@
Let $w\in B$ be fixed. There is a unique bounded solution; all other solutions are oscillating and diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{+\infty}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
@@
$\psi_{w}(\mu(w),t)=-(1-w)\,t^{w}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
From the Hypothesis (H) we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
@@
Now, prove the last item of the present Lemma. From equation (3), we have
$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w)t^{{w}}\int_{t}^{+\infty}u^{-1-w}(\eta(u)-\rho)du,\quad\forall t\geq 1.$
Using the integration by parts formula, we obtain
$\psi_{{w}}(z_{{w}},t)=-\rho\dfrac{1-w}{w}-(1-w^{2})t^{w}\int_{t}^{+\infty}u^{-2-w}\int_{1}^{u}(\eta(v)-\rho)dvdu.$
For every $t\geq 1$, we get
$\left|\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\right|\leq\,t^{-1}\dfrac{|1-w^{2}|}{1+\Re(w)}\sup_{t\geq u}\Big|\int_{1}^{u}(\eta(v)-\rho)dv\Big|.$
By the Hypothesis (H), we obtain
$\sup_{t\geq 1}\left|t\Big(\psi_{{w}}(z_{{w}},t)+\rho\dfrac{1-w}{w}\Big)\right|<+\infty.$
The following theorem, show in particular, that if $\Re(s)\neq\frac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
Theorem 4.
Consider the function $\mu$ given by the Notation 2. For every $s,s^{\prime}\in B$ with $\Re(s)\neq\Re(s^{\prime})$ and $\Im(s)=\Im(s^{\prime})$, we have
$\big(\mu(s),\,\mu(s^{\prime}))\neq(1,1).$
In particular, for every $s\in B$ we have the following implication:
$\big(\mu(s),\,\mu(1-\overline{s})\big)=(1,1)\implies\Re(s)=\frac{1}{2}.$
2 Main Proposition
@@
Let be $s:=\sigma+i\tau\in B$ and $s^{\prime}:=\sigma^{\prime}+i\tau\in B$. Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{s^{\prime}}(1,\cdot)$ of the differential equation (1) as given in Notation 1. Suppose that $\sigma>\sigma^{\prime}$, then
$\Re\big(t^{-i\,\tau}\psi_{s}(1,t)\big)>\Re\big(t^{-i\,\tau}\psi_{s^{\prime}}(1,t)\big),\qquad\forall t>1.$
Proof.
Let be $w\in B$ and denote $w:=\alpha+i\beta$ . For every $t\geq 1$ define
$\psi_{w}(t):=t^{-i\beta}\psi_{w}(1,t).$
According to Notation 1, these functions satisfy the differential equations
$\dot{\psi}_{w}(t)=\alpha\,t^{-1}\psi_{w}(t)+(1-w)\,t^{-1-i\beta}\eta(t),\quad\psi_{w}(1)=1.$
In order to simplify the notation, denote
$x_{w}(t):=\Re\big(\psi_{w}(t)\big)-1,\qquad\forall t\geq 1.$
The previous differential equations, implies that
$\dot{x}_{w}(t)=\alpha\,t^{-1}x_{w}(t)+\alpha t^{-1}+t^{-1}\Re\big((1-w)\,t^{-i\beta}\,\eta(t)\big),\quad x_{w}(1)=0.$ (4)
@@
Let be $s:=\sigma+i\tau\in B$ and $s:=\sigma^{\prime}+i\tau\in B$ such that $\sigma>\sigma^{\prime}$. The previous differential equation implies
$\dot{x}_{s}(t)-\dot{x}_{s^{\prime}}(t)$ $=\sigma\,t^{-1}x_{s}(t)-\sigma^{\prime}\,t^{-1}x_{s^{\prime}}(t)$ (5)
$+(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\,t^{-i\tau}\,\eta(t)\big)\big).$
We have $x_{s}(1)=x_{s^{\prime}}(1)=0$. The previous differential equation implies that
$\dot{x}_{s}(1)-\dot{x}_{s^{\prime}}(1)=(\sigma-\sigma^{\prime})t^{-1}\big(1-\Re\big(\eta(t)\big)\big).$
@@
By the hypothesis (H), we have $|\eta(t)|<1$ for all $t\geq 1$. By hypothesis we have $\sigma>\sigma^{\prime}$, then $\dot{x}_{s}(1)>\dot{x}_{s^{\prime}}(1)$. By continuity there exists $\epsilon>0$ such that
$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,1+\epsilon).$
@@
$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ x_{s}(u)>x_{s^{\prime}}(u)\}.$ (6)
The strategy to prove the present Proposition is to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
$x_{s}(t_{*})=x_{s^{\prime}}(t_{*}),$
@@
$x_{s}(t)>x_{s^{\prime}}(t),\quad\forall t\in(1,t_{*}).$ (7)
We have two cases: $x_{s}(t_{*})\geq 0$ and $x_{s}(t_{*})<0$.
If $x_{s}(t_{*})\geq 0$, then $x_{s^{\prime}}(t_{*})\geq 0$. Using the fact that $\sigma>\sigma^{\prime}$, from the differential equations (5), we get
$\dot{x}_{s}(t_{*})-\dot{x}_{s^{\prime}}(t_{*})$ $=(\sigma-\sigma^{\prime})t_{*}^{-1}x_{s}(t_{*})$
$+(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big)$
$\geq(\sigma-\sigma^{\prime})t_{*}^{-1}\big(1-\Re\big(\,t_{*}^{-i\tau}\,\eta(t_{*})\big)\big).$
By the hypothesis (H), we obtain $\dot{x}_{s}(t_{*})>\dot{x}_{s^{\prime}}(t_{*})$. There exists $t<t_{*}$ close enough to $t_{*}$ such that $x_{s}(t)<x_{s^{\prime}}(t)$. We have obtained a contradiction with equation (7).
If $x_{s}(t_{*})<0$, then $x_{s^{\prime}}(t_{*})<0$. By continuity there exists $\tilde{\epsilon}>0$ such that
$0>x_{s}(u)>x_{s^{\prime}}(u),\quad\forall u\in(t_{*}-\tilde{\epsilon},t_{*}).$

### v4

@@ Abstract
In this manuscript, we introduce a family of parametrized nonhomogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip. We identify a suitable rigidity hypothesis under which a phenomenon of structural asymmetry is established: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$.
@@ 1 Main results
In this manuscript, we study the following non-homogeneous linear complex parametrized differential equations
$\dot{x}=\Re(w)\,t^{-1}x+(1-w)\,t^{-1-i\,\Im(w)}\eta(t),$ (1)
@@
where $w\in\mathbb{C}$ and $\eta:[1,+\infty)\to\mathbb{C}$ is a locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$. The equation (1) can be viewed as a differential equation defined on $1,+\infty)$ with values in $\mathbb{R}^{2}$, and we are interested in the initial conditions of the bounded solutions. This result provides an answer to the Dynamical Conjecture formulated in [2]. In the study that follows, we fix a specific function $\eta$ which is locally integrable function and satisfy $|\eta(t)|<1$ for all $t\geq 1$. The function
$t\mapsto t^{\Re(w)}\int_{1}^{t}u^{-1-w}\eta(u)\,du,\quad t\geq 1,$
@@
$\psi_{w}(z,t)=t^{\Re(w)}\Big[z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big],\quad\forall t\geq 1.$ (2)
@@
For every $w\in\mathbb{C}$ and $z\in\mathbb{C}$, we denote by $\psi_{w}(z,\cdot):[1,+\infty)\to\mathbb{C}$ the unique continuous solution of the differential equation (1) given by equation (2). The point $z$ is called the initial condition of the solution $\psi_{w}(z,\cdot)$.
@@
Denote by $B\subset\mathbb{C}$ the following set
$B:=\Big\{w\in\mathbb{C}:\quad\Re(w)>0\Big\}.$
Let $\mu$ denote the function $\mu:B\mapsto\mathbb{C}$, defined as
@@
where the function $\mu$ is given by the Notation 2.
@@
Let $w\in B$ be fixed. There is a unique bounded solution; all other solutions diverge in norm. Let $z\in\mathbb{C}$ and suppose that $\sup_{t\geq 1}|\psi_{w}(z,t)|<+\infty$. By equation (2), this implies
$\sup_{t\geq 1}\Bigg[t^{\Re(w)}\Big|z+(1-w)\int_{1}^{t}u^{-1-w}\eta(u)\,du\Big|\Bigg]<+\infty.$
@@
$\psi_{w}(\mu(w),t)=-(1-w)\,t^{\Re(w)}\int_{t}^{+\infty}u^{-1-w}\eta(u)\,du,\quad\forall t\geq 1.$ (3)
By Hypothesis we have $\sup_{t\geq 1}|\eta(t)|\leq 1$, we obtain
@@
Let be $s:\sigma+i\tau\in B$. Consider the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) as given in Notation 1. The function $\delta_{s}$ defined as
$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
satisfies the following two equalities:
$\dot{\delta}_{s}(1)$ $=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big),$
${\delta}_{s}(t)$ $=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$\ +4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right),\qquad\forall t\geq 1.$
According to the notation 1 consider the following rigidity hypothesis H:
  • H.
    For every $\beta\in\mathbb{R}$ and every $t>1$,
    $\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right)>0.$
The previous hypothesis suggest the rigidity of the sign. It is suitable in order to apply the results of this dynamical study to the Riemann zeta function.
The following theorem shows that, under the rigidity hypothesis H, if $\Re(s)\neq\tfrac{1}{2}$, then the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ of the differential equation (1) cannot both be bounded on $1,+\infty)$.
Theorem 5.
Consider the function $\mu$ defined in Notation 2. Under hypothesis H, for every $s\in\mathbb{C}$ with $\Re(s)\in(0,1)/\{\frac{1}{2}\}$, we have
$\big(\mu(s),\,\mu(1-\overline{s})\big)\neq(1,1).$
2 Proof of the Proposition 4 and Theorem 5
Proof of the Proposition 4.
Let be $w\in B$ and denote $w:=\alpha+i\beta$. For every $t\geq 1$ denote
$x_{w}(t):=t^{-\frac{1}{2}}\Re\big(\psi_{w}(1,t)\big).$ (4)
According to Notation 1, we obtain the following differential equation
$\dot{x}_{w}(t)=\big(\alpha-\frac{1}{2}\big)\,t^{-1}x_{w}(t)+t^{-\frac{3}{2}}\Re\big((1-w)\,t^{-i\beta}\,\eta(t)\big),\quad x_{w}(1)=1.$ (5)
@@ 2 Proof of the Proposition 4 and Theorem 5
Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(\frac{1}{2},1)$. From the previous equation, we have
$\dot{x_{s}}(t)+\dot{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big({x_{s}}(t)-{x}_{1-\bar{s}}(t)\big)+t^{-\frac{3}{2}}\Re\Big((1-2i\tau)t^{-i\tau}\,\eta(t)\Big).$
Use the fact that $x_{s}(1)=x_{1-\bar{s}}(1)=1$ and integrate, we obtain
${x_{s}}(t)+{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
$+2+\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
From the differential equation (5), we have
$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)$ $=(\sigma-\frac{1}{2})t^{-1}\big(x_{s}(t)+x_{1-\bar{s}}(t)\big)$
$-(2\sigma-1)t^{-\frac{3}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big).$
The two previous differential equations, gives
$\dot{x}_{s}(t)-\dot{x}_{1-\bar{s}}(t)=$ $(\sigma-\frac{1}{2})^{2}t^{-1}\int_{1}^{t}u^{-1}\big({x_{s}}(u)-{x}_{1-\bar{s}}(u)\big)\,du$
$+2(\sigma-\frac{1}{2})t^{-1}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
$+(\sigma-\frac{1}{2})t^{-1}\Re\Big((1-2i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
Define
$\delta_{s}(t):=\Re\big({\psi_{s}}(1,t)-{\psi}_{1-\bar{s}}(1,t)\big),\quad\forall t\geq 1.$
By the notation (4), we have ${x_{s}}(t)-{x}_{1-\bar{s}}(t):=t^{-\frac{1}{2}}\delta_{s}(t)$. We obtained
$\dot{\delta}_{s}(t)=$ $\frac{1}{2}t^{-1}\delta_{s}(t)+(\sigma-\frac{1}{2})^{2}t^{-\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}}\delta_{s}(u)\,du$
$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Big(1-t^{-\frac{1}{2}}\Re\big(\,t^{-i\tau}\,\eta(t)\big)\Big)$
$+2(\sigma-\frac{1}{2})t^{-\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{3}{2}-i\tau}\,\eta(u)\,du\Big).$
Since ${\delta}_{s}(1)=0$ and since by hypothesis we have $|\eta(1)|<1$, then
$\dot{\delta}_{s}(1)=(2\sigma-1)\Big(1-\Re\big(\eta(1)\big)\Big).$
Now, prove the second item of the present Proposition. Integrate the previous equation, we obtain
${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
$+2(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\,du\Big).$
Use the integration by pars formula to the last integral, we get
${\delta}_{s}(t)=\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}\dot{\delta}_{s}(v)\,dv\,du$
$+2(\sigma-\frac{1}{2})\int_{1}^{t}u^{-\frac{1}{2}}\Re\Big(1-u^{-\frac{1}{2}-i\tau}\,\eta(u)\Big)du$
$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big)$
$-4(\sigma-\frac{1}{2})\Re\Big((\frac{1}{2}-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big).$
Equivalent to
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})t^{\frac{1}{2}}\big[(\frac{1}{2}-i\tau)\int_{1}^{t}v^{-\frac{3}{2}-i\tau}\,\eta(v)\,dv\Big]$
$-4(\sigma-\frac{1}{2})\Big[1+\Re\Big((1-i\tau)\int_{1}^{t}u^{-1-i\tau}\,\eta(u)\,du\Big)\Big].$
According to Notation 1, we obtain
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\beta}(1,t)-\psi_{i\beta}(1,t)\right).$
Proof of the Theorem 5.
Let be $s:=\sigma+i\tau\in B$ such that $\sigma\in(0,1)/\{\frac{1}{2}\}$. By symmetry, we may assume that $\sigma\in(\frac{1}{2},1)$. For every $t\geq 1$ denote
$\delta_{s}(t):=\Re\big(\psi_{s}(1,t)-\psi_{1-\bar{s}}(1,t)\big).$
@@
Thank to the Proposition 4, we have $\dot{\delta}_{s}(1)>0$. By continuity there exists $\epsilon>0$ such that
$\delta_{s}(t)>0,\quad\forall t\in(1,1+\epsilon).$
@@
$t_{*}:=\sup\{t>1:\forall\ 1<u<t,\ \delta_{s}(u)>0\}.$
To prove of the present Proposition it is equivalent to prove that $t_{*}=+\infty$. By contradiction, suppose that $t_{*}<+\infty$. Then
$\delta_{s}(t_{*})=0,$ (6)
@@
$\delta_{s}(u)>0,\quad\forall u\in(1,t_{*}).$ (7)
Thank to the Proposition 4, we have
${\delta}_{s}(t)=$ $\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du$
$+4(\sigma-\frac{1}{2})\Re\!\left(\psi_{\frac{1}{2}+i\,\tau}(1,t)-\psi_{i\tau}(1,t)\right).$
Since $2\sigma-1>0$, thanks to the hypothesis H, we obtain
${\delta}_{s}(t)>\frac{1}{2}\int_{1}^{t}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$ (8)
We obtain
${\delta}_{s}(t_{*})>\frac{1}{2}\int_{1}^{t_{*}}u^{-1}\delta_{s}(u)du+(\sigma-\frac{1}{2})^{2}\int_{1}^{t_{*}}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,$
Since $2\sigma-1>0$, by equation (7) we get
${\delta}_{s}(t_{*})>0$
We obtained a contradiction with equation (6). Then $t_{*}=+\infty$. In other words, $\delta_{s}(t)>0$ for all $t>1$. The equation (8) implies
$\dot{\delta}_{s}(t)>(\sigma-\frac{1}{2})^{2}\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{2}v^{-\frac{3}{2}}{\delta}_{s}(v)\,dv\,du,\quad\forall t>1.$
That gives
${\delta}_{s}(t)>2(\sigma-\frac{1}{2})^{2}(t^{\frac{1}{2}}-1)\int_{1}^{2}u^{-\frac{3}{2}}{\delta}_{s}(u)du,\quad\forall t>1.$
Then $\lim_{t\to+\infty}\delta_{s}(t)=+\infty$. This implies that the continuous solutions $\psi_{s}(1,\cdot)$ and $\psi_{1-\bar{s}}(1,\cdot)$ are not both bounded on $1,+\infty)$. By Lemma 3, we deduce that
$(1,1)\neq\big(\mu(s),\,\mu(1-\overline{s})\big).$
More precisely, the proof of Proposition 4 shows that under the rigidity hypothesis H, for all $s\in B$ with $\Re(s)\in(\frac{1}{2},1)$, the function $\mu$ satisfies the following equation
$\Re\big(\mu(s)\big)-\Re\big(\mu(1-\overline{s})\big)\leq 2\Re\big(\eta(1)\big).$
In this case, $\dot{\delta}_{s}(1)\leq 0$; That is, the function does not necessary start positively from 11. Hence the proof of the Theorem 5 cannot be continued.
3 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$. We will now see that hypothesis H, is also satisfied. From the equation (2), we have
$\psi_{w}(1,t)=t^{\Re(w)}\Big(1+(1-w)\int_{1}^{t}u^{-1-w}\,\{u\}\,du\Big),\quad\forall w\in\mathbb{C},\quad\forall t>1,$
which can be written as
$\psi_{w}(1,t)$ $=t^{\Re(c)}\Big[1+J(t,\lfloor t\rfloor)+\sum_{n=1}^{\lfloor t\rfloor-1}J(n+1,n)\Big],$
$J_{w}(x,n)$ $:=(1-w)\int_{n}^{x}u^{-1-w}\,(u-n)\,du,$
with $\sum_{n=1}^{\lfloor t\rfloor-1}J_{w}(n+1,n)=0$ when $t\in[1,2)$. Use the integration by parts formula, we obtain
$J_{w}(x,n)$ $=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,\frac{d}{du}\Big(u^{-1}\,(u-n)\Big)du$
$=\int_{n}^{x}\Big(x^{1-w}-u^{1-w}\Big)\,u^{-2}du.$
For every $\beta\in\mathbb{R}$ and $t\geq 2$ we have
$\Big|t^{\frac{1}{2}}J_{\frac{1}{2}+i\,\beta}(x,n)-J_{i\,\beta}(x,n)\Big|$ $\leq\int_{n}^{x}\Big(\big(t^{\frac{1}{2}}x^{\frac{1}{2}}-x\big)+(t^{\frac{1}{2}}u^{\frac{1}{2}}-u)\,\Big)u^{-2}du$